Calculate the work done by a long distance runner during a 10 km race. The mass of the runner is 52kg . Gravitational field strength is9.8N/kg
1
Expert's answer
2016-11-11T09:44:10-0500
W=mgh=52*9.8*10000=5.1*10^6 J.